Harmonic sequence:
2,2/3,2/5,2/7 with solution
What is the next term in the harmonic sequence 1/11,1/15,1/19,1/23,...? With solution
What is the nth term of the harmonic sequence 1/2,1/4,1/6,1/8 with solution

Answer by Edwin McCravy(20077)

Harmonic sequence:

2,2/3,2/5,2/7 with solution 

A harmonic sequence is the sequence of reciprocals of the terms of
an arithmetic sequence.

So we make the corresponding arithmetic sequence by taking reciprocals
of the terms:

1/2, 3/2, 5/2, 7/2, which has a1 = 1/2, and d = 2/2 or 1

If has nth term an=a1+(n-1)d
                an=1/2+(n-1)(1)
                an=1/2+n-1
                an=n-1/2
                an=2n/n-1/n
                an=(2n-1)/n

So the nth term of the original harmonic sequence is the reciprocal of that:

                an=n/(2n-1)

-------------------

For the others, form the corresponding arithmetic sequence, and proceed as 
above.

Edwin